# Learning Causally Disentangled Representations via the Principle of Independent Causal Mechanisms

### v3

@@ 1 Introduction
Related Work. Our work builds upon the ideas presented in iVAE Khemakhem et al. (2020) and causal variants Yang et al. (2021); Komanduri et al. (2022) and extends them to consider a principled view of causal disentanglement in the label supervised setting. DIVA Ilse et al. (2020) and CCVAE Joy et al. (2021) are special case implementations of the iVAE framework. Yang et al. Yang et al. (2021) proposed CausalVAE, which uses a causal masking layer and is limited to linear SCMs. Komanduri et al. Komanduri et al. (2022) extended this to a nonlinear setting and proposed a causal prior. Both works proposed simplistic models to learn causal mechanisms under the strictly additive noise assumption and do not, from an empirical or theoretical perspective, focus on disentanglement. Shen et al. Shen et al. (2022) proposed learning causal representations supervised by a GAN loss. There has also been work focusing on learning causal representations from paired counterfactual data Brehmer et al. (2022), temporal data Lippe et al. (2023), in self-supervised learning von Kügelgen et al. (2021), and when interventional data is available Ahuja et al. (2023). Unlike many previous works in label-supervised VAE-based CRL, we consider general nonlinear SCMs instead of restricting to additive noise models, propose a causal prior to causally factorize the latent space, and achieve disentanglement of causal mechanisms, as summarized in Table 1.
Contributions. (1) We propose ICM-VAE, a framework for causal representation learning under supervision from labels, where causal variables are derived from nonlinear flow-based diffeomorphic causal mechanisms. (2) Based on the ICM principle, we propose the notion of causal disentanglement for causal models from the perspective of mechanisms and design a causal disentanglement prior to causally factorize the learned distribution over causal variables. (3) Utilizing the structure from our causal disentanglement prior, we theoretically show the identifiability of the learned causal factors up to permutation and elementwise reparameterization. (4) We experimentally validate our method and show that our model can almost perfectly disentangle the causal factors, improve interventional robustness, and generate consistent counterfactual instances in the label-supervised setting.
@@ 2.2 Structural Causal Model
where $\forall i$, $f_{i}:\mathcal{E}_{i}\times\prod_{j\in\textbf{pa}_{i}}\mathcal{Z}_{j}\to\mathcal{Z}_{i}$ are causal mechanisms that determine each causal variable as a function of the parents and noise, $z_{\textbf{pa}_{i}}$ are the parents of causal variable $z_{i}$; and a probability measure $p_{\mathcal{E}}(\epsilon)=p_{\mathcal{E}_{1}}(\epsilon_{1})p_{\mathcal{E}_{2}}(\epsilon_{2})\dots p_{\mathcal{E}_{n}}(\epsilon_{n})$, which admits a product distribution. For the purposes of this work, we assume a causally sufficient setting (no hidden confounding), no SCM misspecification, and faithfulness is satisfied. We depict the causal structure of $z$ by a causal directed acyclic graph (DAG) $\mathcal{G}^{z}$ with adjacency matrix $G^{z}\in\{0,1\}^{n\times n}$.
@@ 4.1 Structural Causal Flow
where $\epsilon_{\textbf{pa}_{i}}$ denotes the noise terms associated with the parents of causal variable $z_{i}$.
@@ 5 Identifiability Analysis
We design our framework to satisfy the conditions necessary to achieve causal mechanism equivalence and causally disentangle the factors of variation. The causally factorized prior in Eq. (11) induces disentanglement of causal mechanisms. Theorem 1 extends the identifiability theorem from iVAE Khemakhem et al. (2020) to show causal mechanism equivalence identifiability when we have a causal model. We note that the causal mechanism disentanglement implies the disentanglement of causal factors.
@@
For a detailed proof of Theorem 1, refer to Appendix A.1.
@@ 6 Experimental Evaluation
In this section, we empirically evaluate the effectiveness of ICM-VAE. We consider a setting where the causal variables can be multi-dimensional and fix the dimension of each causal variable, which has also been explored in previous work Yang et al. (2021); Lippe et al. (2023). This assumption enables the representation to learn more informative and specific codes to describe the factors of variation (i.e. x-position, y-position, etc.). We run experiments on datasets with continuous factors, but discrete flows Tran et al. (2019) can be used to learn causal mechanisms for discrete factors. Our results suggest a component-wise correspondence between learned and true causal factors.
@@
Datasets. We show the performance of our framework on three datasets, each consisting of four real-valued causal variables. The Pendulum dataset Yang et al. (2021) consists of causal variables with causal graph (pendulum angle $\to$ shadow length, shadow position) and (light position $\to$ shadow length, shadow position). The Flow dataset Yang et al. (2021) consists of variables with causal graph (ball size $\to$ water height), (hole position $\to$ water flow), and (water height $\to$ water flow). We also show experiments on a more complex 3D dataset of a robot arm interacting with colored buttons called CausalCircuit Brehmer et al. (2022), which consists of variables with causal graph (robot arm $\to$ blue light intensity, green light intensity, and red light intensity), (blue light intensity $\to$ red light intensity), and (green light intensity $\to$ red light intensity).
@@
Evaluation Metrics. The DCI metric Eastwood and Williams (2018) quantifies the degree to which ground-truth factors and learned latents are in one-to-one correspondence. We compute the DCI disentanglement ($D$) and completeness ($C$) scores, which are based on a feature importance matrix quantifying the degree to which each latent code is important for predicting each ground truth causal factor. The informativeness ($I$) score is the prediction error in the latent factors predicting the ground-truth generative factors and is constant ($I=0$) throughout all datasets and models, so we omit it for brevity. We train models with $3$ random seeds and select the median DCI score to report. We note that DCI is highly correlated with other disentanglement metrics, such as MCC, with strong connections to identifiability Eastwood et al. (2023). To evaluate how changes in the generative factors affect the latent factors, we compute the interventional robustness score (IRS) Suter et al. (2019), which is similar to an $R^{2}$ value.
@@
Implementation. For the Pendulum ($6$K training and $1$K testing) and Flow ($6$K training and $2$K testing) datasets, we linearly increase the $\beta$ parameter throughout training from 00 to $1$. We train for $9\cdot 10^{3}$ steps using a batch size of $64$. We use a Gaussian encoder and decoder with mean and variance computed by fully connected neural networks. For the CausalCircuit dataset ($35$K training and $10$K testing), we linearly increase $\beta$ from 00 to $0.05$. We train for $3.5\cdot 10^{4}$ steps using a batch size of $100$. We use a convolutional neural network architecture with $6$ layers and ReLU activation followed by a fully connected layer to estimate the mean and variance. The noise level for the variance of the Gaussian distribution of the conditional prior is controlled by $\sigma^{2}\in\{0.01,0.00001\}$. The structural causal flow and $\bm{\lambda}$ are implemented as affine form autoregressive flows with the slope and offset computed by fully connected 3-layer neural networks with $100$ unit hidden layers and ReLU activation. We set the learning rate to $0.001$ for all experiments. We set the dimension of each causal variable to $4$ for all datasets. Our code is available at https://github.com/Akomand/ICM-VAE.
@@
Baselines. We compare the performance of our approach, in terms of disentanglement and interventional robustness, with four baseline models: $\beta$-VAE Higgins et al. (2017) (unsupervised and acausal), iVAE Khemakhem et al. (2020) (acausal), CausalVAE Yang et al. (2021) (causal), and SCM-VAE Komanduri et al. (2022) (causal). The causal baselines propose relatively simplistic models that do not necessarily guarantee the disentanglement of causal factors.
@@
Causal Disentanglement. Our experiments show that learning diffeomorphic causal mechanisms, rather than linear SCM, and incorporating the causal structure to learn a bijective $\bm{\lambda}$ to estimate the parameters of the causally factorized distribution significantly improves the disentanglement and interventional robustness of learned causal factors compared with baselines, as shown in Table 2. Consistent with our intuition, iVAE fails to disentangle the causal factors. The results indicate that ICM-VAE disentangles the causal factors and mechanisms almost perfectly. A high DCI disentanglement score indicates a permutation matrix mapping the latent factors to ground-truth generative factors in an ideal one-to-one mapping Eastwood et al. (2023). Further, our model improves the interventional robustness of the representation, where interventions on ground-truth factors map to interventions on the corresponding learned factors.
Table 2: Causal Disentanglement of ICM-VAE and baselines
@@
Counterfactual Generation. We show counterfactual generated results of intervening on learned latent factors. Figure 4 shows the CausalCircuit system and the result of intervening on the robot arm factor and propagating causal effects. We observe that the red light also turns on as the robot arm interacts with the blue or green lights. On the other hand, when the arm interacts with the red light, only the red light turns on and the other lights remain off. We observe a similar phenomenon in the Pendulum and Water Flow systems in Figure 3, which shows the result of intervening on causal factors and propagating effects. Intervening on the pendulum angle or light position has causal effects on the shadow. However, interventions on the shadow factors do not change the parent factors. For the counterfactual generation procedure and results from intervening on other causal factors and iVAE latent traversals, refer to Appendix B.
@@ 7 Conclusion
In this work, we propose a framework for causal representation learning under supervision from labels. We model causal mechanisms as learnable flow-based diffeomorphic transformations from noise to causal variables. We propose the notion of causal mechanism disentanglement for causal models and a causal disentanglement prior, which causally factorizes the learned distribution over causal variables. We theoretically show permutation-equivalent identifiability of the learned factors. We experimentally validate our method and show that our model almost perfectly disentangles the causal factors, improves interventional robustness, and generates consistent counterfactual instances. Future work includes exploring disentanglement given only partially observed labels and leveraging the structure of Jacobians for causal discovery Reizinger et al. (2023).
8 Acknowledgements
@@ A.2 Traditional disentanglement cannot guarantee independent causal mechanisms equivalence
In this section, we provide a counterexample to show that the traditional notion of disentanglement cannot capture the equivalence of causal mechanisms. For example, consider the running example of the Pendulum system. We have four causal factors that are causally related. Let $z$ denote the true factors of variation and $\hat{z}$ denote the learned factors, where each $z_{i}$ and $\hat{z}_{i}$ correspond to the same causal variable. For the sake of simplicity, suppose we have a linear SCM:
Ground-truth SCM
[…]
Learned SCM
[…]
Marginal
[…]
Observe that the causal mechanisms learned are different than the true causal mechanisms. In the linear case, for simplicity, we swap the coefficients. We have the following equivalent marginal distribution for true and learned factors:
@@
However, traditional disentanglement does not imply an equivalence of all the individual causal mechanisms of the true and learned factors. In the above example, the true SCM consists of different mechanisms than the learned SCM, but both yield the same marginal distribution. This example violates the causal mechanism permutation equivalence and causal disentanglement but satisfies traditional disentanglement. We claim that learning a model that achieves equivalence of causal mechanisms from the perspective of the ICM principle better captures disentanglement in the causal setting.
@@ B.1 Dataset Details
Causal Circuit. The Causal Circuit dataset is a new dataset created by Brehmer et al. [2022] to explore research in causal representation learning. The dataset consists of $512\times 512\times 3$ resolution images generated by $4$ ground-truth latent causal variables: robot arm position, red light intensity, green light intensity, and blue light intensity. The images show a robot arm interacting with a system of buttons and lights. The data is rendered using an open-source physics engine. The original dataset consists of pairs of images before and after an intervention has taken place. For the purposes of this work, we only utilize observational data of either the before or after system. The data is generated according to the following process:
@@ B.5 Baselines
We note that none of the identifiability results from baselines focus on causal mechanism equivalence. Our proposed prior encourages a causally factorized latent space, which induces a mechanism equivalence and causal disentanglement.

### v4

@@ 1 Introduction
Related Work.
Our work builds upon the ideas presented in iVAE Khemakhem et al. (2020) and causal variants Yang et al. (2021); Komanduri et al. (2022) and extends them to consider a principled view of causal disentanglement in the label supervised setting. DIVA Ilse et al. (2020) and CCVAE Joy et al. (2021) are special case implementations of the iVAE framework. Yang et al. Yang et al. (2021) proposed CausalVAE, which uses a causal masking layer and is limited to linear SCMs. Komanduri et al. Komanduri et al. (2022) extended this to a nonlinear setting and proposed a causal prior. Both works proposed simplistic models to learn causal mechanisms under the strictly additive noise assumption and do not, from an empirical or theoretical perspective, focus on disentanglement. Shen et al. Shen et al. (2022) proposed learning causal representations supervised by a GAN loss. There has also been work focusing on learning causal representations from paired counterfactual data Brehmer et al. (2022), temporal data Lippe et al. (2023), in self-supervised learning von Kügelgen et al. (2021), and when interventional data is available Ahuja et al. (2023). Unlike many previous works in label-supervised VAE-based CRL, we consider general nonlinear SCMs instead of restricting to additive noise models, propose a causal prior to causally factorize the latent space, and achieve disentanglement of causal mechanisms, as summarized in Table 1.
Contributions.
(1) We propose ICM-VAE, a framework for causal representation learning under label supervision, where causal variables are derived from nonlinear flow-based diffeomorphic causal mechanisms. (2) Based on the ICM principle, we propose the notion of causal disentanglement for causal models from the perspective of mechanisms and design a causal disentanglement prior to causally factorize the learned distribution over causal variables. (3) Using the structure from our causal disentanglement prior, we theoretically show identifiability of the learned causal factors up to permutation and elementwise reparameterization. (4) We experimentally validate our method and show that our model can almost perfectly disentangle the causal factors, improve interventional robustness, and generate consistent counterfactuals.
@@ 2.2 Structural Causal Model
where $\forall i$, $f_{i}:\mathcal{E}_{i}\times\prod_{j\in\textbf{pa}_{i}}\mathcal{Z}_{j}\to\mathcal{Z}_{i}$ are causal mechanisms that determine each causal variable as a function of the parents and noise, $z_{\textbf{pa}_{i}}$ are the parents of causal variable $z_{i}$; and a probability measure $p_{\mathcal{E}}(\epsilon)=p_{\mathcal{E}_{1}}(\epsilon_{1})p_{\mathcal{E}_{2}}(\epsilon_{2})\dots p_{\mathcal{E}_{n}}(\epsilon_{n})$, which admits a product distribution. For the purposes of this work, we assume a causally sufficient setting (no hidden confounding), no SCM misspecification, and faithfulness is satisfied. We depict the causal structure of $z$ by a directed acyclic graph (DAG) $\mathcal{G}^{z}$ with adjacency matrix $G^{z}\in\{0,1\}^{n\times n}$.
@@ 4.1 Structural Causal Flow
where $\epsilon_{\textbf{pa}_{i}}$ denotes the noise terms associated with the parents of causal variable $z_{i}$. The structural causal flow can easily be generalized to multivariate scenarios by masking groups of latent codes corresponding to each causal variable.
@@ 5 Identifiability Analysis
We design our framework to satisfy the conditions necessary to achieve causal mechanism equivalence and causally disentangle the factors of variation. The causally factorized prior in Eq. (11) induces disentanglement of causal mechanisms. Theorem 1 extends the identifiability theorem from iVAE Khemakhem et al. (2020) to show causal mechanism equivalence identifiability when we have a causal model. We note that causal mechanism disentanglement implies the disentanglement of causal factors.
@@
For the proof of Theorem 1, refer to Appendix A.1.
@@ 6 Experimental Evaluation
In this section, we empirically evaluate the effectiveness of ICM-VAE. We consider a setting where the causal variables can be multi-dimensional and fix the dimension of each causal variable, which has also been explored in previous work Yang et al. (2021); Lippe et al. (2023). This assumption enables the representation to learn more informative and specific latent codes to describe the factors of variation (i.e. x-position, y-position, etc.). We run experiments on datasets with continuous factors, but discrete flows Tran et al. (2019) can be used for discrete factors. Our results suggest a component-wise correspondence between the learned and true causal factors.
Datasets.
@@
We show the performance of our framework on three datasets, each consisting of four real-valued causal variables. The Pendulum dataset Yang et al. (2021) consists of causal variables with causal graph (pendulum angle $\to$ shadow length, shadow position) and (light position $\to$ shadow length, shadow position). The Flow dataset Yang et al. (2021) consists of variables with causal graph (ball size $\to$ water height), (hole position $\to$ water flow), and (water height $\to$ water flow). We also show experiments on a more complex 3D dataset of a robot arm interacting with colored buttons called CausalCircuit Brehmer et al. (2022), which consists of variables with causal graph (robot arm $\to$ blue light intensity, green light intensity, and red light intensity), (blue light intensity $\to$ red light intensity), and (green light intensity $\to$ red light intensity).
Evaluation Metrics.
@@
The DCI metric Eastwood and Williams (2018) quantifies the degree to which ground-truth factors and learned latents are in one-to-one correspondence. We compute the DCI disentanglement ($D$) and completeness ($C$) scores, which are based on a feature importance matrix quantifying the degree to which each latent code is important for predicting each ground truth causal factor. The informativeness ($I$) score is the prediction error in the latent factors predicting the ground-truth generative factors and is constant ($I=0$) throughout all datasets and models, so we omit it for brevity. We train models with $3$ random seeds and select the median DCI score to report. We note that DCI is highly correlated with other disentanglement metrics, such as MCC, with strong connections to identifiability Eastwood et al. (2023). To evaluate how changes in the generative factors affect the latent factors, we compute the interventional robustness score (IRS) Suter et al. (2019), which is similar to an $R^{2}$ value.
Implementation.
@@
For the Pendulum ($6$K training and $1$K testing) and Flow ($6$K training and $2$K testing) datasets, we linearly increase the $\beta$ parameter throughout training from 00 to $1$. We train for $9\cdot 10^{3}$ steps using a batch size of $64$. We use a Gaussian encoder and decoder with mean and variance computed by fully connected neural networks. For the CausalCircuit dataset ($35$K training and $10$K testing), we linearly increase $\beta$ from 00 to $0.05$. We train for $3.5\cdot 10^{4}$ steps using a batch size of $100$. We use a convolutional neural network architecture with $6$ layers and ReLU activation followed by a fully connected layer to estimate the mean and variance. The noise level for the variance of the Gaussian distribution of the conditional prior is controlled by $\sigma^{2}\in\{0.01,0.00001\}$. The structural causal flow and $\bm{\lambda}$ are implemented as affine form autoregressive flows with the slope and offset computed by fully connected 3-layer neural networks with $100$ unit hidden layers and ReLU activation. We set the learning rate to $0.001$ for all experiments. We set the dimension of each causal variable to $4$ for all datasets. Our code is available at https://github.com/Akomand/ICM-VAE.
Baselines.
@@
We compare the performance of our approach, in terms of disentanglement and interventional robustness, with four baseline models: $\beta$-VAE Higgins et al. (2017) (unsupervised and acausal), iVAE Khemakhem et al. (2020) (acausal), CausalVAE Yang et al. (2021) (causal), and SCM-VAE Komanduri et al. (2022) (causal). The causal baselines propose relatively simplistic models that do not necessarily guarantee the disentanglement of causal factors.
Causal Disentanglement.
@@
Our experiments show that learning diffeomorphic causal mechanisms, rather than linear SCM, and incorporating the causal structure to learn a bijective $\bm{\lambda}$ to estimate the parameters of the causally factorized distribution significantly improves the disentanglement and interventional robustness of learned causal factors compared with baselines, as shown in Table 2. Consistent with our intuition, iVAE fails to disentangle the causal factors. The results indicate that ICM-VAE disentangles the causal factors and mechanisms almost perfectly. A high DCI disentanglement score indicates a permutation matrix mapping the latent factors to ground-truth generative factors in an ideal one-to-one mapping Eastwood et al. (2023). Further, our model improves the interventional robustness of the representation, where interventions on ground-truth factors map to interventions on the corresponding learned factors.
@@
Table 2: Causal Disentanglement of ICM-VAE and baselines
Counterfactual Generation.
We show counterfactual generated results of intervening on learned latent factors. Figure 4 shows the CausalCircuit system and the result of intervening on the robot arm factor and propagating causal effects. We observe that the red light also turns on as the robot arm interacts with the blue or green lights. On the other hand, when the arm interacts with the red light, only the red light turns on and the other lights remain off. We observe a similar phenomenon in the Pendulum and Water Flow systems in Figure 3, which shows the result of intervening on causal factors and propagating effects. Intervening on the pendulum angle or light position has causal effects on the shadow. However, interventions on the shadow factors do not change the parent factors. For the counterfactual generation procedure, results from intervening on other causal factors, and iVAE latent traversals, refer to Appendix B.3.
@@ 7 Conclusion
We propose ICM-VAE, a framework for causal representation learning under label supervision. We model causal mechanisms as flow-based transformations from noise to causal variables. We extend the idea of disentanglement to causal models and propose the notion of causal mechanism disentanglement. To this end, we design a causal disentanglement prior to causally factorize the distribution over causal variables. We theoretically show permutation-equivalent identifiability of the learned factors. Experimental results show that ICM-VAE almost perfectly disentangles the causal factors, improves interventional robustness, and generates consistent counterfactuals. Future work will incorporate causal discovery and disentanglement given partially observed labels.
Acknowledgements
@@ A.2 Traditional disentanglement cannot guarantee independent causal mechanisms equivalence
In this section, we provide a counterexample to show that the traditional notion of disentanglement cannot capture the equivalence of causal mechanisms. For example, consider the running example of the Pendulum system. We have four causal factors that are causally related. Let $z$ denote the true factors of variation and $\hat{z}$ denote the learned factors, where each $z_{i}$ and $\hat{z}_{i}$ correspond to the same causal variable. For the sake of simplicity, consider the example in Figure 5.
Observe that the causal mechanisms learned are different than the true causal mechanisms. We have the following equivalent marginal distribution for true and learned factors:
@@
However, traditional disentanglement does not imply equivalence of all individual causal mechanisms of true and learned factors. In the above example, the true SCM consists of different mechanisms than the learned SCM, but both yield the same marginal distribution. This example violates the causal mechanism permutation equivalence and causal disentanglement but satisfies traditional disentanglement. We claim that learning a model that achieves equivalence of causal mechanisms from the perspective of the ICM principle better captures disentanglement in the causal setting.
@@ B.1 Dataset Details
Causal Circuit. The Causal Circuit dataset is a new dataset created by Brehmer et al. [2022] to explore research in causal representation learning. The dataset consists of $512\times 512\times 3$ resolution images generated by $4$ ground-truth latent causal variables: robot arm position, red light intensity, green light intensity, and blue light intensity. The images show a robot arm interacting with a system of buttons and lights. The data is rendered using an open-source physics engine. The original dataset consists of pairs of images before and after an intervention has taken place. For this work, we only utilize observational data of either the before or after system. The data is generated according to the following process: